OpenStudy (anonymous):

Can someone check if i did this right? (-9x +5)^1/2 = (3 + 4x)^1/2

OpenStudy (anonymous):

The ^1/2 are like if you are going to square root it afterwards but anyway

OpenStudy (anonymous):

First step combine like terms (-5x^2 + 8) = 0

OpenStudy (anonymous):

Then -1(5x - 8)^1/2 = 0

OpenStudy (anonymous):

-1(5x)^1/2 = 8

OpenStudy (anonymous):

Opps i ment -1(5x^2)^1/2 = 8

OpenStudy (anonymous):

-1(x^2)^1/2 = 8/5

OpenStudy (anonymous):

\[-x = -\frac{ 8 }{ 5 }\]

OpenStudy (anonymous):

And i think that is it.. Is this correct?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Yay thx